How much fence will I need to buy for a yard of 1800 square feet?
Answer by gc2000(22)

Normally more info is given than in this question. Regardless, we can say a few things:
The perimeter of a rectangle decreases the closer the shape [of a constant square footage (in this case 1800)] is to a square; the perimeter increases the more elongated it is.


So these extremes can be given:
(assuming the yard is not less than 1ft wide)
Most fencing: a 1ft x 1800ft (1800SF) yard could require 1800 + 1800 + 1 + 1 = 3602 linear feet of fencing.


Least fencing:
a square yard, (sides each 30 times the square root of 2) would require 30*4*SR2 = "120 times the square root of 2, linear feet = about 169.71 ft" of fencing


to calculate the SR (square root) of 1800, factor out squares:
1800 = 10 * 10 * 3 * 3 * 2
the 10's and 3's factor out, but the 2 doesn't.
one side = 10 * 3 *(SR2) = 30*SR2 (approx. 42.43 ft)
30*SR2 times four equal sides =
exactly 120*SR2 ft.
= approx. 169.71 ft.